This document has a little bit of text, and a little bit of equations. First, we can put really simple mathematical expressions like π or sin(x) into the text by surrounding

them with dollar signs. We can put simple equations into the text in the same way, like this: y = 2x.

But if we want to display a complicated mathematical expression, we really ought to place it on a line all by itself. Surround the equation with pairs of dollar signs, like this:

y = A sin(ωt) + B cos(ωt)